OpenStudy (anonymous):

What is the sin(-330)?

hartnn (hartnn):

sin(-330)=sin(360-330)=sin 30 u know whats sin 30?

OpenStudy (anonymous):

1/2

hartnn (hartnn):

right,so sin(-330) is also 1/2

OpenStudy (anonymous):

What about cos and tan?

hartnn (hartnn):

same way,cos 30 and tan 30 because -330 is same as 30

OpenStudy (anonymous):

So which one would be negative?

hartnn (hartnn):

none of them -330 or 30 lies in 1st quadrant where all ratios,sin,cos,tan,cot......are positive

OpenStudy (anonymous):

What about -240?

hartnn (hartnn):

-240=360-240=120 this lies in 2nd quadrant only sin and cosec will be positive